OpenStudy (dukelaxfan):

The pool company developed new chemicals that transform the pH scale. Using the pH function p(t) = −log10t as the parent function, explain which transformation results in a y-intercept and why. You may graph by hand or using technology. Use complete sentences and show all translations on your graph. p(t) + 1 p(t + 1) −1 • p(t)

OpenStudy (dukelaxfan):

@ganeshie8 @mathmale @whpalmer4 @mathstudent55

OpenStudy (dukelaxfan):

If you ever swam in a pool and your eyes began to sting and turn red, you felt the effects of an incorrect pH level. pH measures the concentration of hydronium ions and can be modeled by the function p(t) = −log10t. The variable t represents the amount of hydronium ions; p(t) gives the resulting pH level. Water at 25 degrees Celsius has a pH of 7. Anything that has a pH less than 7 is called acidic, a pH above 7 is basic, or alkaline. Seawater has a pH just more than 8, whereas lemonade has a pH of approximately 3.

OpenStudy (dukelaxfan):

I've tried graphing on GeoGebra, but it wont graph unless I create sliders

jhonyy9 (jhonyy9):

p(t)= -log10t so than you know this function how you will get y ?

OpenStudy (dukelaxfan):

switch p(t) to y

jhonyy9 (jhonyy9):

yes right so than you know that y = -log10t what is the base of log in this case ?

OpenStudy (dukelaxfan):

10

OpenStudy (dukelaxfan):

\[y=\log_{10}(t) \]

jhonyy9 (jhonyy9):

so than y = - lg(t) yes ?

OpenStudy (dukelaxfan):

what happened to the 10 and the o

jhonyy9 (jhonyy9):

so try calculi p(t) +1 = ? p(t+1) = ? and -1*p(t) = ? what will get ?

OpenStudy (dukelaxfan):

i dont understand what goes in place of question mark

jhonyy9 (jhonyy9):

why ? these all are inside the text of this exercise - yes ? so you have wrote there above

OpenStudy (dukelaxfan):

This is my weakest topic of algebra 2, and i dont get it that well.

jhonyy9 (jhonyy9):

so for example when you need graph a function f(x)= x +2 how you get and how you know what will be the x an y value what you need marking on the graph using axes x and y ? so just use these knowledge in this case of exercise too

jhonyy9 (jhonyy9):

p(t) will be y and t sign x - yes ?

OpenStudy (dukelaxfan):

sure

jhonyy9 (jhonyy9):

so than just use them here now

OpenStudy (dukelaxfan):

so y+1=?

jhonyy9 (jhonyy9):

y=?

OpenStudy (dukelaxfan):

i still dont get it

jhonyy9 (jhonyy9):

just assum to y the 1

jhonyy9 (jhonyy9):

p(t) = - lg t yes ? so p(t)+1 = -lg t +1 but you can rewriting 1 = lg 10 so than p(t)+1 = -lg t + lg 10 can you continue it ?

OpenStudy (dukelaxfan):

no, you are explaining it well, i just dont get it at all

jhonyy9 (jhonyy9):

so you know that lg a + lg b = lg (a*b) yes ?

OpenStudy (dukelaxfan):

yes

jhonyy9 (jhonyy9):

- lg t = lg t^(-1) = lg 1/t --- so is this true ?

OpenStudy (dukelaxfan):

ye

jhonyy9 (jhonyy9):

p(t)+1 = -lg t + lg 10 = lg 1/t + lg 10 = lg 10/t - right ?

OpenStudy (dukelaxfan):

i guess

jhonyy9 (jhonyy9):

so now try calculi it the second case - ok ?

OpenStudy (dukelaxfan):

p(t+1)=?

jhonyy9 (jhonyy9):

yes

OpenStudy (dukelaxfan):

p(t+1)=-log(t+1)

jhonyy9 (jhonyy9):

so and with minus in front of log what you make ?

OpenStudy (dukelaxfan):

log(t+1)^-1

jhonyy9 (jhonyy9):

p(t)+1 = lg 10/t try get it in this style

jhonyy9 (jhonyy9):

so than p(t+1) = ?

OpenStudy (dukelaxfan):

ill try

OpenStudy (dukelaxfan):

log10*t

jhonyy9 (jhonyy9):

how you have got it so ?

OpenStudy (mathmale):

@dukelaxfan: Is this the problem in which you want help? If no, please say specifically what you need to know / learn.

OpenStudy (dukelaxfan):

learn how to get the answers

OpenStudy (dukelaxfan):

i dont get what jhonyy it trying to teach me

jhonyy9 (jhonyy9):

ok. @mathmale please try it you

jhonyy9 (jhonyy9):

@DukeLaxFan you dont wan to graph these points for you can get the tranformations of y ?

OpenStudy (mathmale):

Duke, again I ask YOU to explain in words what you want to know.

OpenStudy (dukelaxfan):

Yes, I need to graph them

jhonyy9 (jhonyy9):

@DukeLaxFan you have wrote this above - yes ? explain which transformation results in a y-intercept and why. You may graph by hand

OpenStudy (dukelaxfan):

or technology

OpenStudy (dukelaxfan):

i want to know how Using the pH function p(t) = −log10t as the parent function, explain which transformation results in a y-intercept and why

OpenStudy (dukelaxfan):

and i can graph using technology if its relevant

OpenStudy (mathmale):

OK. Graph the parent function, p(t) = −log10t. Use the Draw utility, below.

OpenStudy (dukelaxfan):

can i just put a picture of it

OpenStudy (mathmale):

You could use your graphing calculator for this and then copy the graph to the Draw utility, or take and upload a photo of your calculator screen.

OpenStudy (mathmale):

Very nice graph, appropriate. Now, suppose you take the parent function p(t) and re-write it by adding 6. What happens to the graph? Use the online graphing utility if possible to show this.

OpenStudy (dukelaxfan):

sarcasm?.

OpenStudy (mathmale):

Absolutely not! What makes you think I'm being sarcastic?

OpenStudy (dukelaxfan):

just making sure

OpenStudy (mathmale):

Let's concentrate on the math problem at hand.

OpenStudy (dukelaxfan):

yessir

OpenStudy (dukelaxfan):

so p(t)+6=-log(10)(t)

OpenStudy (mathmale):

Actually, it'd be -log t + 6. Use a different name: p2 (t) = -log t + 6 since you are changing the original function, p(t).

OpenStudy (dukelaxfan):

do i include the base aswell?

OpenStudy (mathmale):

No. p(t) = −log10t. p2(t)=-log t + 6. Graph this new function. Note: I am not ignoring your (10) when you write log(10)t. "log t" implies base 10, so it's not necessary to write (10).

OpenStudy (dukelaxfan):

oh ok, i didnt know

OpenStudy (mathmale):

So, graph this new function p2(t) as you did p(t), EXCEPT now add 6 to the function (right side ONLY). Let's see your graph.

OpenStudy (dukelaxfan):

i couldnt do the p2 for some reason

OpenStudy (dukelaxfan):

My wifi is acting up

OpenStudy (dukelaxfan):

hope that works

OpenStudy (mathmale):

Your graph of p2(t) is fine. Does it have a y-intercept?

OpenStudy (dukelaxfan):

i dont think so, if it does, ittle take me an hour to find

OpenStudy (mathmale):

the whole point is to determine whether or not each version of the original graph has a y-intercept. p(t) and p2(t) do NOT. right?

OpenStudy (dukelaxfan):

no

OpenStudy (mathmale):

Right. Now type out p(t) again. Next, throw out the 't' and replace it with (t+1).

OpenStudy (dukelaxfan):

it says i need sliders

OpenStudy (mathmale):

??? p(t) = -log t Throw out the t: p( ) = - log ( ) Write "t+1" where the "t" was.

OpenStudy (dukelaxfan):

sorry zeronknight

OpenStudy (dukelaxfan):

test

OpenStudy (mathmale):

Good morning! Next move: graph p(t+1). Do you recognize this as a form of p(t-a) or p(t+b)? In the case of p(t-a), move your entire graph "a" units to the RIGHT. In the case of p(t+b), move your entire graph "b" units to the LEFT. This is called "horizontal translation."

OpenStudy (dukelaxfan):

Same to you. Yes I remember the horizontal and vertical shifts

OpenStudy (mathmale):

So, you have this lovely graph of p(t)=-log t. Will you be translating this graph to left or to the right as you portray p(t+1)?

OpenStudy (dukelaxfan):

left

OpenStudy (mathmale):

Right. We translate the graph to the left. Graph this new function p(t+1). Does the graph intersect the y-axis? If so, at what point (approx.)?

OpenStudy (dukelaxfan):

p(t+1)=-log(t)?

OpenStudy (dukelaxfan):

anytime i try to graph it, it says to create sliders

OpenStudy (dukelaxfan):

Im gonna be afk for a few